OpenStudy (anonymous):

find inverse of y=-8x+16 i don't need answer please just explain how to solve

Nnesha (nnesha):

replace x with a y and y with a x variable

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

how did you get ?? how did you get 6 at the numerator ?

OpenStudy (anonymous):

I meant 16 not 6 sorry i got y= x-16/-8

Nnesha (nnesha):

alright simplify

OpenStudy (anonymous):

y=x+2

Nnesha (nnesha):

uh-oh

Nnesha (nnesha):

\[\frac{ x-16}{ -8 } ~~means~~ \frac{ x }{ -8 } -\frac{ 16 }{ -8 }\] both terms are dividing by -8

OpenStudy (anonymous):

ok so it would be x/-8 - 2

Nnesha (nnesha):

yes you can write it as \[-\frac{ 1 }{ 8 }x ~~or~~ x\frac{ x }{ 8 }\]

Nnesha (nnesha):

wait no -16/-8 = ?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha yes you can write it as \[-\frac{ 1 }{ 8 }x ~~or~~ x\frac{ x }{ 8 }\] \(\color{blue}{\text{End of Quote}}\) sorry typo \[-\frac{ 1 }{ 8 }x ~~or~~ -\frac{ x }{ 8 }\]

OpenStudy (anonymous):

is it so its -1/8x+2

Nnesha (nnesha):

yes right +2

OpenStudy (anonymous):

thank you

Nnesha (nnesha):

my pleasure :=)